OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3).

OpenStudy (anonymous):

y = -1/2x + 6

OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x-3y=-12 and passes through the point (-3,4).

OpenStudy (anonymous):

4x-3y=-24

OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x-3y=-12 and passes through the points (-3,4). Answer: y=4/3x+8 y=4/3x+3 y=-3/4x+8 y=-3/4x+3

OpenStudy (anonymous):

y=4/3x+8

OpenStudy (anonymous):

Which of the following is the slope between the points (0,-10) and (-4,0)

OpenStudy (anonymous):

-5/2

OpenStudy (anonymous):

Which of the following is the slope between the points (-1,-5) and (7,-3)

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation x-4y=4

OpenStudy (anonymous):

x intercept : (4, 0) y intercept: (0, -1)

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2,-5) with a slope of -3

OpenStudy (anonymous):

y = -3x + 1

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (3,-2) and parallel to y=3x-6

OpenStudy (anonymous):

y=3x-11

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2,6) and parallel to y=-3x-4

OpenStudy (anonymous):

y = -3x+12

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2,4) and perpendicular to y=-1/2x+4

OpenStudy (anonymous):

y = 2x+8

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (4,-2) and perpendicular to y=1/3x-7

OpenStudy (anonymous):

y = 3x-14

OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of 2 and a y-intercept of -3.

OpenStudy (anonymous):

y = 2x-3

OpenStudy (anonymous):

Which of the following is the slope between the two points (4,2) and (3,1)

OpenStudy (anonymous):

solve the compound inequality 6b < 42 or 4b+12 > 8